OpenStudy (anonymous):

what is arctan(sqrt3)?

OpenStudy (campbell_st):

here is a triangle that may help 

OpenStudy (anonymous):

not helpful

OpenStudy (campbell_st):

ok.... what do you know about the tan ratio..?

OpenStudy (anonymous):

that it is opposite/adjacent

OpenStudy (campbell_st):

ok... so you want \[\frac{opp}{adj} = \frac{\sqrt{3}}{1}\] which angle are you looking at?

OpenStudy (anonymous):

yeah but the answer is pi/3 how do u get that?

OpenStudy (campbell_st):

lol... ok... so what is pi/3 in degrees..... pi = 180..... so pi/3 = ? and if you need an answer in radians... please include that information in the question.  you are being asked about an exact value angle... it's worthwhile knowing the exact value triangles